# On a vanishing theorem for surfaces

### v1

@@ 3. Lemmas
Let $f\colon X\to Y$ be a bimeromorphic morphism from a smooth complex analytic surface $X$, and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $X$ such that $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Assume that $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef over $Y$.
@@
Since $f$ is bimeromorphic, $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef and $f$-big over $Y$. By definition, $C\to f(C)$ is finite for any irreducible component $C$ of $\lfloor\Delta-B\rfloor$. Therefore, Lemma 3.1 implies
@@ Acknowledgments.
The first author was partially supported by JSPS KAKENHI Grant Numbers JP21H04994, JP23K20787.

### v2

@@ 3. Lemmas
Let $f\colon X\to Y$ be a projective bimeromorphic morphism from a smooth complex analytic surface $X$, and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $X$ such that $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Assume that $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef over $Y$.
@@
Since $f$ is projective bimeromorphic, $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef and $f$-big over $Y$. By definition, $C\to f(C)$ is finite for any irreducible component $C$ of $\lfloor\Delta-B\rfloor$. Therefore, Lemma 3.1 implies
@@ Acknowledgments.
The first author was partially supported by JSPS KAKENHI Grant Numbers JP21H04994 and JP23K20787. The authors are grateful to Yohei Hada for carefully reading the manuscript and kindly pointing out an error in an earlier version of this paper.